• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 14:44:32 on May 09, 2008

    Entry number 237849

    This entry is a followup to: 237848


    keyword=Left Arm Dipole NMR studies done

    Finished with left arm NMR studies. Here's a summary:
    5/9/08 (left-nmr-tests.txt - JJL)
    Tests of Left NMR
    Po set = 1 GeV/c (was 0)
    Probe A quickly gets lock and settles in.

    11:57 switch to probe B (no change- good)

    12:02 switch to probe D
    Field changes at NMR but P0 remains correct,
    magnet current and Hall field stay the same (good!)

    12:07 switch to probe E (again good)

    12:12 back to A
    current=332.47 (1.7e-04 change)

    12:16 back to B
    Above summarized in plot (probe-switchABDE.JPG)

    12:26 back to A
    Meter doesn't go into "search mode". Control panel says it is,
    but it isn't.
    Pushing search OFF then search ON doesn't help.
    Changing the DAC# doesn't help.
    Going to manual mode, tuning the DAQ# to the correct value,
    then going to AUTO doesn't work (positive and negative polarities)

    Probe A does go into search mode but gets hung up in the 0.26-0.28 T

    13:12 going down to zero

    13:21 P0=0.65 Probe A has problems (stuck in the same band)
    13:26 switch to probe D. It goes into search mode and finds the lock.
    13:31 P0=0.75, controls switch to probe A, same problem
    13:33 switch to probe D. Locks in ~one minute.

    The AC Power Switch for the Left Arm is hareboot17, port 8. Try this.
    No help.
    Reboot left arm controls. Still the same.
    Probe D works fine.

    Switch back and forth between A and D a few times at 0.751 GeV/c.
    D works fine. A does not.

    There's something wrong with probe A, possibilities include:

    Simply a bad connection easily fixed by unplugging it from the
    multiplexor box and plugging it back in. (This is not really
    easy. The multiplexor is inside the dipole.)

    Could be a bad channel in the mutliplexor. There are two spare channels,
    We could use a diffent channel. (Again access is a problem.)

    The probe itself could be bad. (We're still inside the magnet)


    Long term solution: Investigate/repair the above possible problems during
    the long summer down.

    Short term solution: use probe D in place of probe A. The cross calibration
    is OK at better than 2E-04. Will carefully analze the data from the strip
    charts to get a better evaluation the 2E-04. D is a better choice than B
    because B won't go as low as D (D is identical to A but on the opposite side
    of the gap).
    Javier will try to have the controls changed to default to D rather than A,
    Making the problem transparent to the user. Users must, however, be
    aware of the change if they want to interpret B as measured with
    the NMR as a value of P0.

    A copy of this log entry has been emailed to: gomez@jlab.org

    Figure 1